2. Four charges. g_{x} = 5.00 * 10 degrees * C q-3.00 x C. q 3 =-2.00*10^ C,q 3 =6.00*10^ C, are situated at the corners of a square of side 4.00m. Find the potential at the center of the square?


1
Expert's answer
2022-02-20T08:20:29-0500

"\vec E=\sum \vec E_i=\frac{2k\sum q\vec r}{r^3},"

"E=3.4~\frac{MN}C."